OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x - y = -6 5x - 3y = -13

OpenStudy (anonymous):

hellppp!

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@RadEn

OpenStudy (texaschic101):

15x - y = -6 --- subtract 15x from both sides -y = -15x - 6 ---multiply by -1 to make x positive y = 15x + 6 now you would sub 15x + 6 in for y in the other equation

OpenStudy (anonymous):

ok so, 5x - 3 * 15x + 6 = -13

OpenStudy (texaschic101):

5x - 3y = -13 5x - 3(15x + 6) = -13

OpenStudy (texaschic101):

You need to put parenthesis....that way you know to distribute the -3 through the entire parenthesis

OpenStudy (texaschic101):

understand ? Do you have any questions ?

OpenStudy (anonymous):

so i got this 5x - 45x + 18 = 13

OpenStudy (texaschic101):

check again....remember you are multiplying everything in the parenthesis by a -3. And don't forget that your 13 is negative.

OpenStudy (anonymous):

oh haha, ok let me work this out on paper, my head isnt kinda foggy atm

OpenStudy (anonymous):

40x = 5 is what i came out with?

OpenStudy (anonymous):

ok i got the answer, thanks so much.

OpenStudy (anonymous):

Solve the following system of equations. x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21 (2, 3, 5) (-2, 3, 5) (2, -3, 5) (2, 3, -5)

OpenStudy (anonymous):

is there a formula to solve this.

OpenStudy (texaschic101):

5x - 3(15x + 6) = -13 5x - 45x - 18 = -13 --- add 18 to both sides 5x - 45x = -13 + 18 -- combine like terms -40x = 5 -- divid by -40 x = 5/-40 x = -1/8 now you would sub -1/8 in for x in either equation to find y

OpenStudy (texaschic101):

you would have to use elimination or substitution to solve a system of equations with 3 variables. I myself, prefer elimination. But thats just me.

OpenStudy (anonymous):

The answer to the first problem was asking what the equation would look like so it was : 5x - 3(15x + 6) = -13

OpenStudy (texaschic101):

yes...that is correct

OpenStudy (anonymous):

so could you set me up with the first step in the 3 equations to find it, they are horrid in my brain, as i have ADD.

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

this can be a little hard to understand.....you take 2 equations, any two. x + 3y - z = 2 --->(-1) x - 2y + 3z = 7 --------------- -x - 3y + z = -2 (result of multiplying by -1) x - 2y + 3z = 7 --------------add -5y + 4z = 5 now take one of the equations you used and the equation you did not use, and eliminate what you eliminated in the first part of the problem x - 2y + 3z = 7 --->(-1) x + 2y - 5z = -21 ----------------- -x + 2y - 3z = -7 (result of multiplying by -1) x + 2y - 5z = -21 -----------------add 4y - 8z = -28 now take your results.. -5y + 4z = 5 --->(2) 4y - 8z = -28 --------------- -10y + 8z = 10 (result of multiplying by 2) 4y - 8z = -28 --------------add -6y = - 18 y = -18/-6 y = 3 now sub 3 in for y in the equation that x was eliminated from -5y + 4z = 5 -5(3) + 4z = 5 -15 + 4z = 5 4z = 5 + 15 4z = 20 z = 5 now sub 3 in for y and 5 in for z in any original equation x + 3y - z = 2 x + 3(3) - 5 = 2 x + 9 - 5 = 2 x + 4 = 2 x = 2 - 4 x = -2 check... x - 2y + 3z = 7 -2 - 2(3) + 3(5) = 7 -2 - 6 + 15 = 7 -8 + 15 = 7 7 = 7 (correct) solution is (-2,3,5)